# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

### v3

@@ 8.2 Proofs in Section 3
$\geq-\operatorname{Tr}\left(xx^{\top}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+(1-\gamma)\left(f_{K}\left(\Sigma\right)-f_{K}\left(\Sigma^{\prime}\right)\right),$ (47)
with equality when $K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}=K^{\prime}$ holds. Let $\{x_{t}^{*}\}_{t=0}^{\infty}$ and $\{u_{t}^{*}\}_{t=0}^{\infty}$ be the state and control sequences generated under $\pi^{*}(\cdot|x)=\mathcal{N}(-K^{*}x,\Sigma^{*})$ with noise sequence $\{w_{t}^{*}\}_{t=0}^{\infty}$ with mean 00 and covariance $W$. Apply Lemma 8.1 and (47) with $\pi$ and $\pi^{*}$ to get:
@@
$\leq\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+f_{K}(\Sigma^{*})-f_{K}(\Sigma),$ (48)
where $f_{K}$ is defined in (16). To analyze the first term in (48), note that
$\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{\sigma_{\min}(R)}\operatorname{Tr}(E_{K}^{\top}E_{K})$ (49)
@@
where the last equation follows from (18).
To analyze the second terms in (48), note that $f_{K}$ is a concave function, thus we can find its maximizer $\Sigma_{K}^{*}$ by taking the gradient of $f_{K}$ and setting it to 00, i.e., $\Sigma_{K}^{*}=\frac{\tau}{2}(R+\gamma B^{\top}P_{K}B)^{-1}.$ Thus,
$f_{K}(\Sigma^{*})-f_{K}(\Sigma)\leq f_{K}(\Sigma_{K}^{*})-f_{K}(\Sigma)\overset{(a)}{\leq}\operatorname{Tr}\left(\nabla f_{K}(\Sigma)^{\top}(\Sigma_{K}^{*}-\Sigma)\right)$ (50)
@@
where $(a)$ follows from the first order concavity condition for $f_{K}$ and $(b)$ is from $0\prec\Sigma\preceq I.$
For the lower bound, consider $K^{\prime}=K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}$ and $\Sigma^{\prime}=\Sigma$ where equality holds in (47). Let $\{x_{t}^{\prime}\}_{t=0}^{\infty}$, $\{u_{t}^{\prime}\}_{t=0}^{\infty}$ be the sequence generated with $K^{\prime},\Sigma^{\prime}$. By $C(K^{*},\Sigma^{*})\leq C(K^{\prime},\Sigma^{\prime}),$ we have
$C(K,\Sigma)-C(K^{*},\Sigma^{*})\geq C(K,\Sigma)-C(K^{\prime},\Sigma^{\prime})=-\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{\prime},u_{t}^{\prime})\right]$
@@
Lemma 8.2 (Smoothness of $f_{K}$ (16)).
Let $K\in\mathbb{R}^{k\times n}$ be given and let $f_{K}$ be defined in (16). Fix $0<a\leq 1$. For any symmetric positive definite matrices $X\in\mathbb{R}^{k\times k}$ and $Y\in\mathbb{R}^{k\times k}$ satisfying $aI\preceq X\preceq I$ and $aI\preceq Y\preceq I$, the following inequality holds
@@
$f_{K}(X)-f_{K}(Y)+\operatorname{Tr}\left(\nabla f_{K}(X)^{\top}(Y-X)\right)$ (51)
@@
Now let us show that there exists a constant $m>0$ (independent of $K$) such that for any $Z\in\mathbb{R}^{k\times k}$ with real positive eigenvalues $a\leq z_{1}\leq\cdots\leq z_{k}$, the following holds:
$-\log(\det(Z))+\operatorname{Tr}(Z-I)\leq m\operatorname{Tr}((Z-I)^{2}).$ (52)
Note that (52) is equivalent to $\sum_{i=1}^{k}-\log(z_{i})+z_{i}-1\leq m\sum_{i=1}^{k}(z_{i}-1)^{2}.$ With elementary algebra, one can verify that when $m:={(-\log\left(a\right)+a-1)}\cdot{\left(a-1\right)^{-2}},$ it holds that $-\log(z)+z-1\leq m(z-1)^{2}$ for all $z\geq a$ and such an $m$ satisfies $m\geq\frac{1}{2}$. Therefore, (52) holds. Combining (51) and (52) with $Z=X^{-1}Y$, we see
$f_{K}(X)-f_{K}(Y)+\operatorname{Tr}\left(\nabla f_{K}(X)^{\top}(Y-X)\right)\leq\frac{\tau m}{2(1-\gamma)}\operatorname{Tr}\left((X^{-1}Y-I)^{2}\right).$
Proof.
(of Lemma 3.3). The first equality immediately results from Lemma 8.1, by taking an expectation with respect to $x$ in (46). The last inequality follows directly from Lemma 8.2. ∎
@@ 8.3 Proofs in Section 4
$aI\preceq\Sigma-\frac{\eta}{1-\gamma}\left(R-\frac{\tau}{2}\Sigma^{-1}+\gamma B^{\top}P_{K}B\right)\preceq I.$ (53)
@@
$\Sigma+\frac{{\eta}\tau}{2(1-\gamma)}\Sigma^{-1}-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B)\succeq\left(a+\frac{{\eta}\tau}{2(1-\gamma)a}\right)I-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B)$
$\succeq\left(a+\frac{\eta}{1-\gamma}\|{R}+\gamma B^{\top}P_{K}B\|\right)I-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B)\succeq aI,$
and
$\Sigma+\frac{{\eta}\tau}{2(1-\gamma)}\Sigma^{-1}-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B)\preceq\left(1+\frac{{\eta}\tau}{2(1-\gamma)}\right)I-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B)$
$\preceq\left(1+\frac{\eta}{1-\gamma}\sigma_{\min}(R)\right)I-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B)\preceq I,$
hence (53).
Second, one can show $aI\preceq\Sigma^{\prime}\preceq I:$ observe that (53) is equivalent to
$aI-\Sigma\preceq-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1})\preceq I-\Sigma.$
Then by multiplying $\Sigma$ to both sides then adding a $\Sigma$ to each term, we have
$a\Sigma^{2}-\Sigma^{3}+\Sigma\preceq\Sigma-\frac{\eta}{1-\gamma}\Sigma(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1})\Sigma\preceq\Sigma^{2}-\Sigma^{3}+\Sigma.$
@@
8.3.2 Proof of Lemma 4.4
For ease of exposition, write $S=S_{K,\Sigma},$ $S^{\prime}=S_{K^{\prime},\Sigma^{\prime}}$, and $S^{*}=S_{K^{*},\Sigma^{*}}$. Let $M_{a}$ be defined in the same way as in Lemma 3.3. Let $f_{K}$ be defined as (16). Then Lemma 3.3 implies
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)$ $=\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)\right)$ (54)
@@
$\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)\right)+2\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}E_{K}\right)$ (55)
@@
where (a) follows from $\eta_{1}\leq\frac{1}{2\|R+\gamma B^{\top}P_{K}B\|}$ and (b) follows from (49).
By Lemma 4.3, $aI\preceq\Sigma^{\prime}\preceq I.$ Using Lemma 8.2, one can derive the following inequality:
@@
$\leq\frac{1}{1-\gamma}\operatorname{Tr}\left(\left(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1}\right)(\Sigma^{\prime}-\Sigma)\right)+M_{a}\operatorname{Tr}\left((\Sigma^{-1}\Sigma^{\prime}-I)^{2}\right)$
$=-\frac{\eta_{2}}{(1-\gamma)^{2}}\operatorname{Tr}\left(\left(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1}\right)\Sigma\left(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1}\right)\Sigma\right)$
$\quad\quad\quad\quad+\frac{\eta^{2}_{2}}{(1-\gamma)^{2}}M_{a}\operatorname{Tr}\left(\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)^{2}\right)$
$\overset{(b)}{\leq}$ $-\frac{\eta_{2}}{2(1-\gamma)^{2}}\operatorname{Tr}\left(\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)^{2}\right).$
Here $(b)$ follows from the inequality
$\eta_{2}=\frac{2(1-\gamma)a^{2}}{\tau}\leq\frac{2(1-\gamma)}{\tau}\left(\frac{\tau}{2\|R+\gamma B^{\top}P_{K}B\|}\right)^{2}\overset{(c)}{\leq}\frac{2(1-\gamma)}{\tau}\overset{(d)}{\leq}\frac{1}{2M_{a}}.$
$(c)$ is obtained from the fact that $\tau\leq 2\sigma_{\min}(R)\leq 2\|R+\gamma B^{\top}P_{K}B\|$, and $(d)$ follows from Lemma 8.2. Meanwhile, observe from (50) that
$f_{K}(\Sigma^{*})-f_{K}(\Sigma)$
$\leq\frac{1}{(1-\gamma)}\operatorname{Tr}\left(\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)\Sigma^{-1}(R+\gamma B^{\top}P_{K}B)^{-1}\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)\right)$
@@
$f_{K}(\Sigma)-f_{K}(\Sigma^{\prime})$ $\leq-\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\left(f_{K}(\Sigma)-f_{K}(\Sigma^{*})\right).$ (56)
Finally, with $\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S^{*}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}$, plug (55) and (56) in (54) to get
$C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)$
$\leq-2\eta_{1}\mu\frac{\sigma_{\min}(R)}{\|S^{*}\|}\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)-\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\left(f_{K}(\Sigma)-f_{K}(\Sigma^{*})\right)$
$\leq-\zeta\left(\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+f_{K}(\Sigma)-f_{K}(\Sigma^{*})\right)$
$\leq-\zeta(C(K,\Sigma)-C(K^{*},\Sigma^{*}))$
where the last inequality follows from (48). Adding $C(K,\Sigma)-C(K^{*},\Sigma^{*})$ to both sides of the above inequality finishes the proof.
@@
$\overset{(a)}{\geq}\frac{1}{1-\gamma}\left(\frac{\tau k}{2}-\frac{\tau}{2}\left(k+k\log(2\pi)\right)-\frac{\tau k}{2}\log(\frac{\tau}{2\sigma_{\min}(R)})\right)$
$=\frac{\tau k}{2(1-\gamma)}\log(\frac{\sigma_{\min}(R)}{\pi\tau})=:M_{\tau},$
where $(a)$ follows from the fact that $\sigma_{\min}(R)\operatorname{Tr}(\Sigma)-\frac{\tau}{2}\left(k+k\log(2\pi)-\frac{\tau}{2}\log\operatorname{det}\Sigma\right)$ is a convex function with respect to $\Sigma$ with minimizer $\frac{\tau}{2\sigma_{\min}(R)}I.$ Thus,
$C(K,\Sigma)$ $\geq\mathbb{E}_{x_{0}\sim\mathcal{D}}\left[x_{0}^{\top}P_{K}x_{0}\right]+\frac{\gamma}{1-\gamma}\operatorname{Tr}(WP_{K})+M_{\tau}\geq\left(\mu+\frac{\gamma}{1-\gamma}\sigma_{\min}(W)\right)\|P_{K}\|+M_{\tau}.$

### v4

@@ 9.2 Proofs in Section 3
(9.10) $\geq-\operatorname{Tr}\left(xx^{\top}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+(1-\gamma)\left(f_{K}\left(\Sigma\right)-f_{K}\left(\Sigma^{\prime}\right)\right),$
with equality when $K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}=K^{\prime}$ holds. Let $\{x_{t}^{*}\}_{t=0}^{\infty}$ and $\{u_{t}^{*}\}_{t=0}^{\infty}$ be the state and control sequences generated under $\pi^{*}(\cdot|x)=\mathcal{N}(-K^{*}x,\Sigma^{*})$ with noise sequence $\{w_{t}^{*}\}_{t=0}^{\infty}$ with mean 0 and covariance $W$. Apply Lemma 9.1 and (9.10) with $\pi$ and $\pi^{*}$ to get:
@@
(9.11) $\leq\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+f_{K}(\Sigma^{*})-f_{K}(\Sigma),$
where $f_{K}$ is defined in (3.4). To analyze the first term in (9.11), note that
(9.12) $\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{\sigma_{\min}(R)}\operatorname{Tr}(E_{K}^{\top}E_{K})$
@@
where the last equation follows from (3.6).
To analyze the second terms in (9.11), note that $f_{K}$ is a concave function, thus we can find its maximizer $\Sigma_{K}^{*}$ by taking the gradient of $f_{K}$ and setting it to 0, i.e., $\Sigma_{K}^{*}=\frac{\tau}{2}(R+\gamma B^{\top}P_{K}B)^{-1}.$ Thus,
(9.13) $f_{K}(\Sigma^{*})-f_{K}(\Sigma)\leq f_{K}(\Sigma_{K}^{*})-f_{K}(\Sigma)\overset{(a)}{\leq}\operatorname{Tr}\left(\nabla f_{K}(\Sigma)^{\top}(\Sigma_{K}^{*}-\Sigma)\right)$
@@
where $(a)$ follows from the first order concavity condition for $f_{K}$ and $(b)$ is from $0\prec\Sigma\preceq I.$ Plug (9.12) and (9.13) into (9.11) to get (3.7).
For the lower bound, consider $K^{\prime}=K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}$ and $\Sigma^{\prime}=\Sigma$ where equality holds in (9.10). Let $\{x_{t}^{\prime}\}_{t=0}^{\infty}$, $\{u_{t}^{\prime}\}_{t=0}^{\infty}$ be the sequence generated with $K^{\prime},\Sigma^{\prime}$. By $C(K^{*},\Sigma^{*})\leq C(K^{\prime},\Sigma^{\prime}),$ we have
(9.14) $C(K,\Sigma)-C(K^{*},\Sigma^{*})\geq C(K,\Sigma)-C(K^{\prime},\Sigma^{\prime})=-\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{\prime},u_{t}^{\prime})\right]$
@@
Lemma 9.4 (Smoothness of $f_{K}$ (3.4)).
Let $K\in\mathbb{R}^{k\times n}$ be given and let $f_{K}$ be defined in (3.4). Fix $0<a\leq 1$. For any symmetric positive definite matrices $X\in\mathbb{R}^{k\times k}$ and $Y\in\mathbb{R}^{k\times k}$ satisfying $aI\preceq X\preceq I$ and $aI\preceq Y\preceq I$,
@@
(9.15) $f_{K}(X)-f_{K}(Y)+\operatorname{Tr}\left(\nabla f_{K}(X)^{\top}(Y-X)\right)$
@@
Now let us show that there exists $m\in\mathbb{R}^{+}$ (independent of $K$) such that for any $Z\in\mathbb{R}^{k\times k}$ with real positive eigenvalues $a\leq z_{1}\leq\cdots\leq z_{k}$, the following holds:
(9.16) $\log(\det(Z))+\operatorname{Tr}(Z-I)\leq m\operatorname{Tr}((Z-I)^{2}).$
Note that (9.16) is equivalent to $\sum_{i=1}^{k}-\log(z_{i})+z_{i}-1\leq m\sum_{i=1}^{k}(z_{i}-1)^{2}.$ With elementary algebra, one can verify that when $m:={(-\log\left(a\right)+a-1)}\cdot{\left(a-1\right)^{-2}},$ it holds that $\log(z)+z-1\leq m(z-1)^{2}$ for all $z\geq a$ and such an $m$ satisfies $m\geq\frac{1}{2}$. Therefore, (9.16) holds. Combining (9.15) and (9.16) with $Z=X^{-1}Y$, we see $f_{K}(X)-f_{K}(Y)+\operatorname{Tr}\left(\nabla f_{K}(X)^{\top}(Y-X)\right)\leq\frac{\tau m}{2(1-\gamma)}\operatorname{Tr}\left((X^{-1}Y-I)^{2}\right).$
Proof 9.6.
(of Lemma 3.3). The first equality immediately results from (9.9) in Lemma 9.1. The last inequality follows directly from Lemma 9.4.
@@ 9.3 Proofs in Section 4
(9.17) $aI\preceq\Sigma-\eta(1-\gamma)^{-1}(R-\frac{\tau}{2}\Sigma^{-1}+\gamma B^{\top}P_{K}B)\preceq I.\vskip-2.84526pt$
@@
$\Sigma+\frac{{\eta}\tau}{2(1-\gamma)}\Sigma^{-1}-\frac{\eta(R+\gamma B^{\top}P_{K}B)}{1-\gamma}\succeq\left(a+\frac{{\eta}\tau}{2(1-\gamma)a}\right)I-\frac{\eta(R+\gamma B^{\top}P_{K}B)}{1-\gamma}$
$\succeq\left(a+\frac{\eta\|{R}+\gamma B^{\top}P_{K}B\|}{1-\gamma}\right)I-\frac{\eta(R+\gamma B^{\top}P_{K}B)}{1-\gamma}\succeq aI,\text{ and }$
$\Sigma+\frac{{\eta}\tau}{2(1-\gamma)}\Sigma^{-1}-\frac{\eta(R+\gamma B^{\top}P_{K}B)}{1-\gamma}\preceq\left(1+\frac{{\eta}\tau}{2(1-\gamma)}\right)I-\frac{\eta(R+\gamma B^{\top}P_{K}B)}{1-\gamma}$
[…]
Second, one can show $aI\preceq\Sigma^{\prime}\preceq I:$ observe that (9.17) is equivalent to $aI-\Sigma\preceq-\frac{\eta}{1-\gamma}(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1})\preceq I-\Sigma.$ Then by multiplying $\Sigma$ to both sides then adding a $\Sigma$ to each term, we have
$a\Sigma^{2}-\Sigma^{3}+\Sigma\preceq\Sigma-\frac{\eta}{1-\gamma}\Sigma(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1})\Sigma\preceq\Sigma^{2}-\Sigma^{3}+\Sigma.\vskip-2.84526pt$
@@
9.4 Proof of Lemma 4.6
For ease of exposition, write $S_{K,\Sigma}, S=S$$^{\prime}=S_{K^{\prime},\Sigma^{\prime}}$, and $S^{*}=S_{K^{*},\Sigma^{*}}$. Let $M_{a}$ be defined in the same way as in Lemma 3.3. Let $f_{K}$ be defined as (3.4). Then Lemma 3.3 implies
(9.18) $C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)$ $=\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)\right)$
@@
(9.19) $\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)\right)+2\operatorname{Tr}\left(S^{\prime}(K^{\prime}-K)^{\top}E_{K}\right)$
@@
where $(a)$ follows from $\eta_{1}\leq({2\|R+\gamma B^{\top}P_{K}B\|})^{-1}$ and $(b)$ follows from (9.12).
By Lemma 4.5, $aI\preceq\Sigma^{\prime}\preceq I.$ Then by Lemma 9.4,
@@
$\leq-\frac{\eta_{2}}{(1-\gamma)^{2}}\operatorname{Tr}\left(\left(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1}\right)\Sigma\left(R+\gamma B^{\top}P_{K}B-\frac{\tau}{2}\Sigma^{-1}\right)\Sigma\right)$
$\quad\quad\quad\quad+\frac{(\eta_{2})^{2}M_{a}}{(1-\gamma)^{2}}\operatorname{Tr}\left(\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)^{2}\right)$
$\overset{(c)}{\leq}$ $-\frac{\eta_{2}}{2(1-\gamma)^{2}}\operatorname{Tr}\left(\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)^{2}\right).$
Here $(c)$ follows from the inequality $\eta_{2}=\frac{2(1-\gamma)a^{2}}{\tau}\leq\frac{2(1-\gamma)}{\tau}\left(\frac{\tau}{2\|R+\gamma B^{\top}P_{K}B\|}\right)^{2}\overset{(d)}{\leq}\frac{2(1-\gamma)}{\tau}\overset{(e)}{\leq}\frac{1}{2M_{a}},$ where $(d)$ is obtained from the fact that $\tau\leq 2\sigma_{\min}(R)\leq 2\|R+\gamma B^{\top}P_{K}B\|$, and $(e)$ follows from Lemma 9.4. Meanwhile, observe from (9.13) that
$f_{K}(\Sigma^{*})-f_{K}(\Sigma)$ $\leq\frac{1}{(1-\gamma)}\operatorname{Tr}\Big(\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)\cdot$
$\hskip 71.13188pt\Sigma^{-1}(R+\gamma B^{\top}P_{K}B)^{-1}\left((R+\gamma B^{\top}P_{K}B)\Sigma-\frac{\tau}{2}I\right)\Big)$
@@
(9.20) $f_{K}(\Sigma)-f_{K}(\Sigma^{\prime})$ $\leq-\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\left(f_{K}(\Sigma)-f_{K}(\Sigma^{*})\right).$
Finally, with $\zeta=\min\left\{\frac{2\mu\eta_{1}\sigma_{\min}(R)}{\|S^{*}\|},\frac{\eta_{2}a\sigma_{\min}(R)}{2(1-\gamma)}\right\}$, plug (9.19) and (9.20) in (9.18) to get $C(K^{\prime},\Sigma^{\prime})-C(K,\Sigma)\leq-\zeta\Big(\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+f_{K}(\Sigma)-f_{K}(\Sigma^{*})\Big)$. The proof is finished by applying (9.11) then adding $C(K,\Sigma)-C(K^{*},\Sigma^{*})$ to both sides.
@@
$\overset{(a)}{\geq}\frac{1}{1-\gamma}\left(\frac{\tau k}{2}-\frac{\tau}{2}\left(k+k\log(2\pi)\right)-\frac{\tau k}{2}\log(\frac{\tau}{2\sigma_{\min}(R)})\right)=M_{\tau},$
where $(a)$ follows from the fact that $\sigma_{\min}(R)\operatorname{Tr}(\Sigma)-\frac{\tau}{2}\left(k+k\log(2\pi)-\frac{\tau}{2}\log\operatorname{det}\Sigma\right)$ is a convex function with respect to $\Sigma$ with minimizer $\frac{\tau}{2\sigma_{\min}(R)}I.$ Thus, $C(K,\Sigma)\geq\mathbb{E}_{x_{0}\sim\mathcal{D}}\left[x_{0}^{\top}P_{K}x_{0}\right]+\frac{\gamma\operatorname{Tr}(WP_{K})}{1-\gamma}+M_{\tau}\geq\left(\mu+\frac{\gamma\sigma_{\min}(W)}{1-\gamma}\right)\|P_{K}\|+M_{\tau}.$